OpenStudy (anonymous):

Use polar coordinates to find the limit... \[\lim_{(x,y) \rightarrow (0,0)}(x ^{2}+y ^{2})\ln(x ^{2}+y ^{2})\]

OpenStudy (anonymous):

Converting to polar coordinates I got... \[\lim_{r \rightarrow 0}r ^{2}\ln(r ^{2})\] but ln(0) doesn't exist, so I'm not sure what to do...

OpenStudy (zarkon):

now it is single variable...use L'Hospitals rule

OpenStudy (anonymous):

Wait, but using L'Hopitals rule... 2r*ln(r^2)+2r I would still have the ln(0)

OpenStudy (zarkon):

\[r^2\ln(r^2)=\frac{\ln(r^2)}{1/r^2}\] then use L'Hospitals

OpenStudy (anonymous):

Alright, I got -r^2, so it would be 0, right?

OpenStudy (zarkon):

0 is correct

OpenStudy (anonymous):

Thanks!